paul paints a room in 6 hours. joel paints the room in 3 hours. how long will it take them to paint togather?
Answer by stanbon(57962)

paul paints a room in 6 hours. joel paints the room in 3 hours. how long will it take them to paint together?
--------
Paul's rate: 1/6 job/hr
Joel's rate: 1/3 job/hr
---
Together rate: 1/x job/hr
===============================
Equation:
rate + rate = together rate
1/6 + 1/3 = 1/x
---
x + 2x = 6
3x = 6
x = 2 hrs ( time to do the job together)
==========
Cheers,
Stan H.
==========